# AVCD: Mitigating Hallucinations in Audio-Visual Large Language Models through Contrastive Decoding

### v1

@@ Abstract
Hallucination remains a major challenge in multimodal large language models (MLLMs). To address this, various contrastive decoding (CD) methods have been proposed that contrasts original logits with hallucinated logits generated from perturbed inputs. While CD has shown promise in vision-language models (VLMs), it is not well-suited for AV-LLMs, where hallucinations often emerge from both unimodal and cross-modal combinations involving audio, video, and language. These intricate interactions call for a more adaptive and modality-aware decoding strategy. In this paper, we propose Audio-Visual Contrastive Decoding (AVCD)—a novel, training-free decoding framework designed to model trimodal interactions and suppress modality-induced hallucinations in AV-LLMs. Unlike previous CD methods in VLMs that corrupt a fixed modality, AVCD leverages attention distributions to dynamically identify less dominant modalities and applies attentive masking to generate perturbed output logits. To support CD in a trimodal setting, we also reformulate the original CD framework to jointly handle audio, visual, and textual inputs. Finally, to improve efficiency, we introduce entropy-guided adaptive decoding, which selectively skips unnecessary decoding steps based on the model’s confidence in its predictions. Extensive experiments demonstrate that AVCD consistently outperforms existing decoding methods. Especially, on the AVHBench dataset, it improves accuracy by 6% for VideoLLaMA2 and 11% for video-SALMONN, demonstrating strong robustness and generalizability.
@@ 1 Introduction
We validate the effectiveness of AVCD through extensive experiments across a range of MLLMs, including evaluations involving image liu2024improved , video cheng2024videollama ; lin2023video , and audio-visual inputs cheng2024videollama ; sun2024video . The results consistently demonstrate that AVCD mitigates hallucinations during inference, regardless of the input modality. Specifically, we evaluate AVCD on the AVHBench dataset sung2024avhbench , a benchmark specifically designed to assess hallucinations in audio-visual settings. When applied to two widely used AV-LLMs, AVCD achieves a 6% relative improvement in accuracy on VideoLLaMA2 cheng2024videollama and 11% on video-SALMONN sun2024video . These improvements underscore AVCD’s robustness and its effectiveness in enhancing the reliability of AV-LLMs for trimodal understanding.
@@ 3.2 Audio-Visual Contrastive Decoding
Attentive masking strategy. Once the dominant modality is identified by analyzing the attention distribution of the final query token ($A_{Q_{K}}$), we perform masking on combinations of the less dominant modalities. Specifically, we apply an attentive masking strategy using a threshold defined by the top $P\%$ of the mean stacked $A_{Q_{K}}$ across transformer layers. Tokens in the non-dominant modalities that exceed this threshold are masked out, intentionally suppressing informative signals in those modalities to produce logits biased toward the dominant modality for CD. Unlike existing methods that directly distort inputs leng2024mitigating ; woo2024ritual ; woo2024don , which may introduce undesired noise due to deviations from the trained input distribution, our attentive masking strategy preserves the structure of the learned input space. This design allows AVCD to avoid the noise from directly perturbing inputs, focusing instead on mitigating hallucinations that arise from cross-modal reasoning huo2024self .
@@
Figure 4: KL divergence comparison of original and masked-modality logits. On the AVHBench sung2024avhbench dataset, AVCD shows smaller divergence from the original logit distribution compared to VCD leng2024mitigating .
Based on our derivation, we expect the difference between the original logits and those from corrupted signals to be minimal. To validate the effectiveness of our attentive masking strategy, we compute the KL divergence between the logit distributions from original and corrupted signals using 100 samples from the AVHBench dataset sung2024avhbench . As shown in Figure 4, our method yields a smaller divergence compared to VCD, indicating that our approach introduces less distortion when masking modalities. This result supports the mathematical approximation that allows moving the addition operation outside the logarithm. Accordingly, we can approximate $\text{logit}(\mathbf{y}_{t}|\mathbf{x}^{v},\mathbf{x}^{l})$ as:
@@ 4 Experiments
VideoLLaMA2 cheng2024videollama Base 81.30$\pm$0.09 78.05 3.14$\pm$0.05
VCD leng2024mitigating 77.66$\pm$0.03 62.44 2.87$\pm$0.04
VCD* 81.49$\pm$0.03 79.51 2.98$\pm$0.08
AVCD 81.58$\pm$0.03 81.95 3.31$\pm$0.03
video-SALMONN sun2024video Base 48.50$\pm$0.06 60.00 1.94$\pm$0.05
VCD leng2024mitigating 41.57$\pm$0.08 59.51 1.65$\pm$0.05
VCD* 49.00$\pm$0.11 65.85 2.18$\pm$0.07
AVCD 49.73$\pm$0.06 66.83 2.28$\pm$0.02
@@ 4.1 Experimental Settings
Dataset and evaluation. We evaluate AV-LLMs on MUSIC-AVQA li2022learning , which tests synchronized audio-visual reasoning with question-answering (QA) pairs derived from the MUSIC dataset. We also use AVHBench sung2024avhbench , designed to assess audio-visual hallucinations using adversarial QA pairs. Since audio-video captioning in AVHBench follows a distinct evaluation protocol, we assess it separately. For video-LLMs, we use MSVD-QA xu2016msr , which involves questions about objects, actions, and events in short video clips, using the first 1,000 test examples. We also use ActivityNet-QA yu2019activitynet , a more challenging benchmark requiring reasoning over long videos with complex temporal understanding.
@@ 4.2 Experimental Results
Audio-visual LLMs. In Table 1, we compare three different types of decoding methods using two models, VideoLLaMA2 and video-SALMONN. We use the term Base to refer to the original decoding method. We first extend VCD leng2024mitigating by incorporating the audio modality, following the original CD formulation. This baseline extension, however, leads to performance degradation when applied directly to AV-LLMs. In contrast, VCD*, which integrates audio using our proposed AVCD formulation (Eq. (10)) with dominant modality estimation, improves performance over the base method in most cases. Finally, AVCD, which further introduces a dominance-aware attentive masking strategy in place of the Gaussian noise used in VCD, consistently achieves the best results. This demonstrates the effectiveness of our approach in processing audio-visual inputs. Notably, on AVHBench, a benchmark for audio-visual hallucinations, AVCD improves accuracy by around 6% for VideoLLaMA2 and 11% for video-SALMONN relative to the base method. In addition to simple QA tasks on MUSIC-AVQA and AVHBench, AVCD also excels in audio-video captioning.
@@
VideoLLaMA2 cheng2024videollama Base 74.43$\pm$0.31 47.19$\pm$0.55
VCD leng2024mitigating 71.30$\pm$0.57 45.65$\pm$0.04
AVCD 75.20$\pm$0.42 48.22$\pm$0.04
Video-LLaVA lin2023video Base 70.20$\pm$0.20 47.48$\pm$0.02
VCD leng2024mitigating 71.80$\pm$0.25 47.25$\pm$0.02
AVCD 72.16$\pm$0.24 48.03$\pm$0.14
@@ 4.3 Discussions
Table 3: Ablation on CD with Eq. Eq. (2) and Eq. Eq. (11). AVCD effectively extends existing CD to trimodal configurations.
@@ A V AV
Base $\checkmark$ $\checkmark$ $\checkmark$ 78.05
w/ Eq. 2 $\checkmark$ $\checkmark$ $\checkmark$ 80.49
w/ Eq. 2 $\checkmark$ $\checkmark$ $\checkmark$ 79.51
w/ Eq. 2 $\checkmark$ $\checkmark$ $\checkmark$ 80.49
w/ Eq. 11 $\checkmark$ $\checkmark$ $\checkmark$ 79.51
AVCD $\checkmark$ $\checkmark$ $\checkmark$ 81.95
Revisiting conventional CD. We extend the conventional approach (Eq. (2)) of performing single-instance CD to handle multiple instances from two less dominant modalities simultaneously in AV-LLMs. As shown in rows 2 to 4 of Table 3, the conventional CD (Eq. (2)) shows improved performance compared to base decoding. This demonstrates that the approach is effective for single-instance CD. By adapting this approach for AV-LLMs, where language is dominant, the following formulation is derived.
@@
In the fifth row of Table 3, we report the performance of CD using Eq. (11). Although there is a slight improvement compared to the base model, it is evident that the performance is lower than when using Eq. (2) with either audio or AV masking. This approach over-emphasizes audio-visual contrasts, as audio-induced hallucinations are partially resolved when CD is applied to the AV domain, and video-induced hallucinations are also partially mitigated.
On the other hand, in the sixth row of Table 3, AVCD effectively addresses imbalance by assigning a positive coefficient to each domain’s output, achieving better performance than both Eq. (2)-based and Eq. (11)-based contrastive decoding, highlighting its effectiveness.
@@
Efficient inference via entropy-guided adaptive decoding. In Figure 6, we examine the trade-off between inference speed and accuracy by evaluating VideoLLaMA2 on the AVHBench dataset under varying thresholds ($\tau$) for the EAD strategy. The x-axis represents the average inference time per generated token, while the y-axis reports the corresponding accuracy. As $\tau$ increases, the model becomes less likely to invoke additional forward passes, resulting in faster inference but smaller performance gains.
We observe that accuracy remains stable when the entropy threshold $\tau$ is set between 0.2 and 0.6, but gradually declines as $\tau$ increases to 0.8 and 1.0. At $\tau=1.0$, most additional decoding steps are skipped, resulting in faster inference and accuracy comparable to that of base decoding. In contrast, setting $\tau=0.8$ yields faster inference than VCD (2.25 vs. 2.5) while still improving accuracy over both the base method (78.05% → 80.98%) and VCD. These results show that our EAD strategy successfully balances speed and performance. They also demonstrate that AVCD enables efficient inference, despite requiring multiple full forward passes.
@@
Audio 77.07 -
Vision 75.12 70.70$\pm$0.14
Language 81.95 74.20$\pm$0.14
Adaptive 81.95 75.20$\pm$0.42
Evaluation of dominant modality recognition. To evaluate the effectiveness of our adaptive dominant modality recognition, we compare our method against a fixed strategy that assumes a single modality is always dominant. We use VideoLLaMA2 for the experiments and evaluate on AVHBench for audio-video-text inputs and MSVD-QA for video-text inputs. As shown in Table 4, when using audio-visual inputs, setting language as the dominant modality achieves performance on par with the adaptive strategy. However, fixing audio or vision as dominant leads to a noticeable drop in accuracy. This suggests that when all three modalities are present, the model identifies language as the dominant modality. This observation is supported by our analysis of the final-layer attention distribution (see Figure A.1 in our Supp. D), where 70% of the attention is directed toward language modality.
@@ D.2 Impacts of $\alpha$
To evaluate the performance variation with respect to changes in $\alpha$, we measure the performance while gradually adjusting the value of $\alpha$. As shown in Figure A.1 (a), the dominance between video and audio is relatively balanced, which motivates us to set $\alpha=\alpha^{v}=\alpha^{a}$. We also observe that even when $\alpha$ is adjusted to reflect the slightly higher dominance of the video modality empirically, the performance remains largely unaffected. For instance, when using $\alpha^{v}=2$ and $\alpha^{a}=2.5$, the accuracy on the AVHBench dataset with VideoLLaMA2 remains at 81.95%, identical to the result obtained with the balanced setting $\alpha=\alpha^{v}=\alpha^{a}=2.5$.
@@
D.3 Impacts of masking ratios
We investigate the impact of the masking ratio $P$ in our attentive masking strategy, which determines the proportion of high-attention tokens to be masked. As shown in Table A.2, we experiment with masking ratios of 25%, 50%, 75%, and 100% (i.e., full masking). Among these, a 50% masking ratio yields the highest accuracy. When the ratio is too low (e.g., 25%), the contrast with the original model output is insufficient, limiting the effectiveness of contrastive decoding. Conversely, overly high masking ratios (75% or 100%) cause large deviations from the original logits, leading to greater error in the logarithmic approximation (see Section A.1) and thus degraded performance.

### v2

@@ Abstract
Hallucination remains a major challenge in multimodal large language models (MLLMs). To address this, various contrastive decoding (CD) methods have been proposed that contrasts original logits with hallucinated logits generated from perturbed inputs. While CD has shown promise in vision-language models (VLMs), it is not well-suited for AV-LLMs, where hallucinations often emerge from both unimodal and cross-modal combinations involving audio, video, and language. These intricate interactions call for a more adaptive and modality-aware decoding strategy. In this paper, we propose Audio-Visual Contrastive Decoding (AVCD)—a novel, training-free decoding framework designed to model trimodal interactions and suppress modality-induced hallucinations in AV-LLMs. Unlike previous CD methods in VLMs that corrupt a fixed modality, AVCD leverages attention distributions to dynamically identify less dominant modalities and applies attentive masking to generate perturbed output logits. To support CD in a trimodal setting, we also reformulate the original CD framework to jointly handle audio, visual, and textual inputs. Finally, to improve efficiency, we introduce entropy-guided adaptive decoding, which selectively skips unnecessary decoding steps based on the model’s confidence in its predictions. Extensive experiments demonstrate that AVCD consistently outperforms existing decoding methods. Especially, on the AVHBench dataset, it improves accuracy by 2% for VideoLLaMA2 and 7% for video-SALMONN, demonstrating strong robustness and generalizability. Our code is available at https://github.com/kaistmm/AVCD.
@@ 1 Introduction
We validate the effectiveness of AVCD through extensive experiments across a range of MLLMs, including evaluations involving image liu2024improved , video cheng2024videollama ; lin2023video , and audio-visual inputs cheng2024videollama ; sun2024video . The results consistently demonstrate that AVCD mitigates hallucinations during inference, regardless of the input modality. Specifically, we evaluate AVCD on the AVHBench dataset sung2024avhbench , a benchmark specifically designed to assess hallucinations in audio-visual settings. When applied to two widely used AV-LLMs, AVCD achieves a 2% relative improvement in accuracy on VideoLLaMA2 cheng2024videollama and 7% on video-SALMONN sun2024video . These improvements underscore AVCD’s robustness and its effectiveness in enhancing the reliability of AV-LLMs for trimodal understanding.
@@ 3.2 Audio-Visual Contrastive Decoding
Attentive masking via zeroing out. Once the dominant modality is identified by analyzing the attention distribution of the final query token ($A_{Q_{K}}$), we perform masking on combinations of the less dominant modalities. Specifically, we apply an attentive masking strategy using a threshold defined by the top $P\%$ of the mean stacked $A_{Q_{K}}$ across transformer layers. Tokens in the non-dominant modalities that exceed this threshold are set to zero, intentionally suppressing informative signals in those modalities to produce logits biased toward the dominant modality for CD. Unlike existing methods that directly distort inputs leng2024mitigating ; woo2024ritual ; woo2024don , which may introduce undesired noise due to deviations from the trained input distribution, our attentive masking strategy preserves the structure of the learned input space. This design allows AVCD to avoid the noise from directly perturbing inputs, focusing instead on mitigating hallucinations that arise from cross-modal reasoning huo2024self .
@@
Table 1: Approximation error under the adaptive plausibility constraint. On the AVHBench sung2024avhbench dataset, AVCD produces a smaller deviation from the original logits compared to VCD leng2024mitigating .
Decoding Masked Modality
$downarrow AV\downarrow$
VCD 0.083 0.015 0.073
AVCD (Ours) 0.032 0.015 0.037
Based on our derivation, we expect the difference between the original logits and those from corrupted signals to be minimal. To validate the effectiveness of our attentive masking strategy, we compute the exact approximation error (defined in Supp. A.1) between the logits from original and corrupted signals on 100 samples from the AVHBench dataset sung2024avhbench , after applying the adaptive plausibility constraint. As shown in Table 1, our method produces smaller errors than VCD across multiple masked modality combinations, indicating that it introduces less distortion when masking modalities. This result supports the mathematical approximation that allows moving the addition operation outside the logarithm. Accordingly, we can approximate $\text{logit}(\mathbf{y}_{t}|\mathbf{x}^{v},\mathbf{x}^{l})$ as:
@@ 4 Experiments
[…]
@@ 4.1 Experimental Settings
Dataset and evaluation. We evaluate AV-LLMs on MUSIC-AVQA li2022learning , which tests synchronized audio-visual reasoning with question-answering (QA) pairs derived from the MUSIC dataset. We also use AVHBench sung2024avhbench , designed to assess audio-visual hallucinations using adversarial QA pairs. We treat the entire dataset as the test set and use the initially released portion as the validation set, which is employed to evaluate inference speed as reported in Figure 5. Since audio-video captioning in AVHBench follows a distinct evaluation protocol, we assess it separately. For video-LLMs, we use MSVD-QA xu2016msr , which involves questions about objects, actions, and events in short video clips, using the first 1,000 test examples. We also use ActivityNet-QA yu2019activitynet , a more challenging benchmark requiring reasoning over long videos with complex temporal understanding.
@@ 4.2 Experimental Results
Audio-visual LLMs. In Table 2, we compare three different types of decoding methods using two models, VideoLLaMA2 and video-SALMONN. We use the term Base to refer to the original decoding method. We first extend VCD leng2024mitigating by incorporating the audio modality, following the original CD formulation. This baseline extension, however, leads to performance degradation when applied directly to AV-LLMs. In contrast, VCD*, which integrates audio using our proposed AVCD formulation (Eq. (10)) with dominant modality estimation, improves performance over the VCD method in most cases. Finally, AVCD, which further introduces a dominance-aware attentive masking strategy in place of the Gaussian noise used in VCD, consistently achieves the best results. This demonstrates the effectiveness of our approach in processing audio-visual inputs. Notably, on AVHBench, a benchmark for audio-visual hallucinations, AVCD improves accuracy by around 2% for VideoLLaMA2 and 7% for video-SALMONN relative to the base method. In addition to simple QA tasks on MUSIC-AVQA and AVHBench, AVCD also excels in audio-video captioning.
@@
[…]
@@ 4.3 Discussions
Table 4: Ablation on CD with Eq. (2) and Eq. (11). AVCD effectively extends existing CD to trimodal configurations.
@@ A V AV
Base $\checkmark$ $\checkmark$ $\checkmark$ 70.52
w/ Eq. 2 $\checkmark$ $\checkmark$ $\checkmark$ 71.88
w/ Eq. 2 $\checkmark$ $\checkmark$ $\checkmark$ 70.16
w/ Eq. 2 $\checkmark$ $\checkmark$ $\checkmark$ 72.07
w/ Eq. 11 $\checkmark$ $\checkmark$ $\checkmark$ 70.94
AVCD $\checkmark$ $\checkmark$ $\checkmark$ 72.15
Revisiting conventional CD. We extend the conventional approach (Eq. (2)) of performing single-instance CD to handle multiple instances from two less dominant modalities simultaneously in AV-LLMs. As shown in rows 2 to 4 of Table 4, conventional CD performs comparably or better than base decoding, confirming its effectiveness for single-instance CD. By simply adapting this approach for AV-LLMs, where language is dominant, the following formulation is derived.
@@
In the fifth row of Table 4, we report the performance of CD using Eq. (11). Although there is a slight improvement compared to the base model, it is evident that the performance is lower than when using Eq. (2) with either audio or AV masking. This approach over-emphasizes audio-visual contrasts, as audio-induced hallucinations are partially resolved when CD is applied to the AV domain, and video-induced hallucinations are also partially mitigated. On the other hand, in the sixth row of Table 4, AVCD effectively addresses imbalance by assigning a positive coefficient to each domain’s output, achieving better performance than both Eq. (2)-based and Eq. (11)-based contrastive decoding, highlighting its effectiveness.
@@
Efficient inference via entropy-guided adaptive decoding. In Figure 5, we show the trade-off between inference speed and accuracy for VideoLLaMA2 on the AVHBench validation set under varying EAD thresholds ($\tau$). The x-axis shows average inference time per generated token, and the y-axis indicates accuracy. Higher $\tau$ reduces additional forward passes, speeding up inference but limiting performance gains. Note that FlashAttention dao2023flashattention is not applied, as AVCD relies on full attention weights.
We observe that accuracy remains stable when the entropy threshold $\tau$ is below 0.6, but gradually declines as $\tau$ increases to 0.8 and 1.0. At $\tau=1.0$, most additional decoding steps are skipped, resulting in faster inference and accuracy comparable to that of base decoding without FlashAttention. In contrast, setting $\tau=0.8$ yields faster inference than VCD (2.25 vs. 2.5) while still improving accuracy over both the base method (78.05% → 80.98%) and VCD. These results show that our EAD strategy successfully balances speed and performance. They also demonstrate that AVCD enables efficient inference, despite requiring multiple full forward passes.
@@
Audio 68.67 -
[…]
Evaluation of dominant modality recognition. To evaluate the effectiveness of our adaptive dominant modality recognition, we compare our method against a fixed strategy that assumes a single modality is always dominant. We use VideoLLaMA2 for the experiments and evaluate on AVHBench for audio-video-text inputs and MSVD-QA for video-text inputs. As shown in Table 5, treating language as the dominant modality matches the adaptive strategy, whereas fixing audio or vision leads to a clear drop in accuracy. This suggests that when all three modalities are present, the model identifies language as the dominant modality. This observation is supported by our analysis of the final-layer attention distribution (see Figure A.1 in our Supp. D), where 70% of the attention is directed toward language modality.
@@ D.2 Impacts of $\alpha$
To evaluate the performance variation with respect to changes in $\alpha$, we measure the performance while gradually adjusting the value of $\alpha$. As shown in Figure A.1 (a), the dominance between video and audio is relatively balanced, which motivates us to set $\alpha=\alpha^{v}=\alpha^{a}$. We also observe that even when $\alpha$ is adjusted to reflect the slightly higher dominance of the video modality empirically, the performance remains largely unaffected. For instance, when using $\alpha^{v}=2$ and $\alpha^{a}=2.5$, the accuracy on the AVHBench validation set with VideoLLaMA2 remains at 81.95%, identical to the result obtained with the balanced setting $\alpha=\alpha^{v}=\alpha^{a}=2.5$.
@@
D.3 Evaluation on OmniBench dataset
Following OmniBench [34], we adopt video-SALMONN as the backbone model, for which the original paper reports an overall accuracy of 35.6%. Under the same setting, we reproduce the benchmark and obtain 33.5% accuracy with Base decoding, which serves as our baseline throughout the experiments.
As summarized in Table A.4, AVCD emerges as the only method that consistently surpasses the base model across most evaluation settings. In contrast, alternative approaches frequently fail to improve upon the baseline, often yielding comparable or even degraded results. This consistent superiority underscores AVCD’s robust generalization capability and resilience to dataset variability, demonstrating its effectiveness even on challenging benchmarks such as OmniBench.
Table A.2: Per component analsysis on AVHBench. Results show that each component—dominance-aware masking, trimodal contrastive decoding, and entropy-guided adaptive decoding—plays a complementary role in improving either accuracy or efficiency.
Method Accuracy (%) Inference Speed (sec/token)
Baseline 74.15 4.4
+ Dominance-aware masking 79.02 4.4
   + Eq. (10) 81.95 4.4
    + EAD (Ours) 81.95 3.1
D.4 Per component analysis
We conduct an ablation study on AVHBench validation set using VideoLLaMA2 as the base model in Table A.2. Baseline refers to a contrastive decoding (CD) setup with randomly selected dominant modality and Eq. (11). Dominance-aware masking improves accuracy by approximately +5%. Trimodal CD with Eq. (10), our proposed extension of CD to trimodal alignment, further improves performance by +2.9%. Entropy-guided adaptive decoding (EAD) maintains accuracy while reducing inference latency by over 30%. These results validate the complementary roles of all three AVCD components as each contributes meaningfully to either accuracy or efficiency.
@@ D.4 Per component analysis
D.5 Impacts of masking ratios
We investigate the impact of the masking ratio $P$ in our attentive masking strategy, which determines the proportion of high-attention tokens to be masked. As shown in Table A.3, we experiment with masking ratios of 25%, 50%, 75%, and 100% (i.e., full masking) and evaluate performance on the AVHBench validation set. Among these, a 50% masking ratio yields the highest accuracy. When the ratio is too low (e.g., 25%), the contrast with the original model output is insufficient, limiting the effectiveness of contrastive decoding. Conversely, overly high masking ratios (75% or 100%) cause large deviations from the original logits, leading to greater error in the logarithmic approximation (see Section A.1) and thus degraded performance.
Table A.4: Evaluation on the OmniBench dataset. AVCD demonstrates robust generalization, consistently surpassing the base model. The following abbreviations indicate the evaluation categories: Action: Action & Activity, Story: Story Description, Plot: Plot Inference, Object: Object Identification & Description, Context: Contextual & Environmental, Identity: Identity & Relationship, Text: Text & Symbols, Count: Count & Quantity.
Decoding Overall Action Story Plot Object Context Identity Text Count
Base 33.5 27.1 27.0 24.1 61.1 30.1 46.9 21.4 7.1
VCD 23.3 18.3 17.4 13.1 44.6 26.2 37.5 7.1 14.3
OPERA 30.8 27.9 23.0 20.7 52.1 36.9 34.4 14.3 7.1
VCD* 31.9 27.1 26.1 21.5 57.4 30.5 43.8 21.4 0.0
AVCD 34.5 28.3 28.7 24.5 60.7 30.5 50.0 21.4 7.1
[…]
Category Base AVCD Improvement
[…]
AV Matching 64.4% 71.2% +6.8%
D.6 Analysis on hallucinations
[…]